Accelerator Update for September 8, 2008 - September 10, 2008
Monday Sept 8
The day shift began with Operations monitoring the Tevatron (TeV) store 6406 (82.73E30) and the Recycler (RR) stash (291.9E10), with the Antiproton Source (Pbar) stacking (21.2E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the MiniBooNE, NuMI, and the Meson MTest experiments taking beam.
At 10:39 AM, store 6406 was lost and many systems tripped off. The TeV quenched at sectors D0, D4, and E1. The stack and stash remained intact. Many Front Ends (FE) and the CAMAC crates lost power. Everything was rebooted. The MI power supply interlock loop (MI PSLOOP) went down.
Operations reported that MI resumed ramping at 10:57 AM.
Booster experts reported at 10:57 AM that they were beginning scheduled work on GMPS, MP02, and a power supply (B:SEXSB).
Operations reported at 10:59 AM some odd behavior on the D0 quadrupoles just before the quench.
An RF expert reported that he was working on BRF16 at 10:59 AM.
The Duty Electrician reported at 11:05 AM that he didn’t find anything wrong at the Master Substation.
At 11:07 AM, Operations reported that a MI Lambertson power supply had tripped off. An expert investigated and had it back on about 40 minutes later.
Operations established beam to MiniBooNE at 11:20 AM.
Operations established beam to Pbar, NuMI, and MTest at 11:22 AM.
Operations prepared D0 and CDF for access at 11:33 AM.
Controls experts discovered that a power glitch in the MAC room was the cause of the quench. Operations reported at 11:58 AM that the following systems had been affected by the glitch:
The MAC computer room
At 12:03 PM, the Cryo Coordinator reported that the vacuum had gone bad at D4. He contacted a vacuum expert.
D0 and CDF came out of access at 12:44 PM.
At 1:54 PM, the Pbar Sweeping magnet power supply tripped off. It reset.
At 2:13 PM, a TeV quadrupole power supply (T:QD15) was reporting that it had a blown fuse. A power supply expert investigated and replaced its regulator.
Operations started ramping the TeV at 2:34 PM.
Operations put the TeV through a dry squeeze at 2:50 PM.
Operations put the TeV through a wet squeeze at 3:51 PM.
Operations reported at 4:04 PM that the Recycler had stashed antiprotons three times during the day shift.
At 4:52 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to also trip. Everything reset.
Operations established store 6409 at 6:57 PM with an initial luminosity of 267.97E30.
At 7:04 PM, a TeV RF station (TRF8) tripped off.
At 7:36 PM, the TeV sector E1 wet engine tripped off. Operators were able to turn the engine back on before it had any affect on the E1 cryo system.
Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.
Tuesday Sept 9
The midnight shift began with Operations monitoring store 6409 (131.13E30) and the stash (203.7E10), with the Pbar stacking (29.5E10), and with the MiniBooNE and NuMI experiments taking beam.
Operations established beam to MTest at 4:35 AM.
Operations terminated store 6409 at 7:34 AM.
Operations put the TeV into shot setup at 7:49 AM.
Operations reported at 8 AM that the Recycler had stashed six times during the midnight shift.
An RF expert reported at 8:27 AM that there was nothing wrong with TRF8.
Cryo system experts reported at 8:40 AM that they had completed their work on the TeV sectors E1 and F4 wet engines.
At 9:52 AM, the NuMI beam permit tripped off due to the MI RF system. Everything reset.
Operations established store 6410 at 9:58 AM with an initial luminosity of 271.32E30.
Operations prepared MI-31 and the Pelletron for access at 10:07 AM.
At 10:13 AM, TRF8 tripped off. This trip caused the loss of about 5E30 of luminosity. An RF expert investigated.
Pelletron experts came out of access around 10:47 AM.
At 11:13 AM, the MI PSLOOP tripped off at MI-50. It reset.
The Neutron Therapy Facility reported at 11:25 AM that they had treated their last patient for the day.
Operations prepared the Pelletron for a 20-minute access at 1:15 PM.
Operations prepared the Pelletron for a short access at 3:27 PM.
At 3:34 PM, KRF5 tripped off. It reset.
Operations reported at 4:06 PM that the Recycler had stashed four times during the midnight shift.
At 4:32 PM, KRF7 tripped off. It reset.
Operations disable the MTest beam at 6:07 PM.
At 8:17 PM, a Pbar Debuncher septum (D:ISEPV) tripped off. It reset.
At 10:11 PM, MIRF2 tripped off and caused the NuMI beam permit to also trip off. Everything reset.
At 10:54 PM, BRF8 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed seven times during the evening shift.
Wednesday Sept 10
The midnight shift began with Operations monitoring store 6410 (71.4E30) and the stash (340.5E10), with Pbar stacking (29.6E10), and with the MiniBooNE and NuMI experiments taking beam.
At 1:06 AM, LRF4 tripped off. It reset.
Operations terminated store 6410 at 2:25 AM.
At 2:32 AM, MIRF2 tripped off and took down the NuMI beam permit. Everything reset.
Operations put the TeV into shot setup at 2:37 AM.
At 2:52 AM, the P1 transfer line beam permit tripped off. It reset.
At 4:28 AM, the P1 transfer line beam permit tripped off. Operations reported that an autotune program was driving a power supply (HT701) outside of its limits. Everything reset. Operators restored a good file to the program.
Operations established store 6411 at 5:01 AM with an initial luminosity of 255.37E30.
At 6:14 AM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 and 8 to also trip off. Everything reset.
At 6:45 AM, Operations reported that the Kautz Road substation MOS 86 breaker tripped off. The Duty Electrician investigated and reported that a raccoon had crawled into the harmonic filter. They will have to clean up the raccoon remains before they can turn back on.
At 7:13 AM, KRF7 tripped off. It reset.
At 7:27 AM, a cryo system expert from CHL reported a problem communicating with a compressor at TeV sector C0. He said that its IO crate had a bad power supply.
Operations reported at 7:58 AM that the Recycler had stashed twice during the midnight shift.
Operations prepared the Pelletron for a quick access at 8:01 AM.
The Plans for Wednesday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 09/10/2008 email Fermilab|
|Security, Privacy, Legal|